If 0.200 grams CCl4 is mixed with 0.175 grams of SbF3, how many grams of CCl2F2 is formed


1
Expert's answer
2021-04-15T02:03:59-0400

Solution:

Balanced chemical equation:

3CCl4 + 2SbF3 → 3CCl2F2 + 2SbCl3

According to the equation above: n(CCl4)/3 = n(SbF3)/2 = n(CCl2F2)/3


The molar mass of CCl4 is 153.82 g mol-1.

The molar mass of SbF3 is 178.75 g mol-1.

The molar mass of CCl2F2 is 120.91 g mol-1.


Determine moles of each reactant:

0.200 g CCl4 × (1 mol CCl4 / 153.82 g CCl4) = 0.001300 mol CCl4

0.175 g SbF3 × (1 mol SbF3 / 178.75 g SbF3) = 0.000979 mol SbF3


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it. Let's choose CCl4:

n(SbF3) = 2 × n(CCl4) / 3 = (2 × 0.001300 mol) / 3 = 0.000867 mol

 

The calculation above means that we need 0.000867 mol of SbF3 to completely react with CCl4.

We have 0.000979 mol of SbF3 and therefore more than enough antimony trifluoride.

Thus antimony trifluoride (SbF3) is in excess and carbon tetrachloride (CCl4) is the limiting reagent.


Hence,

n(CCl4)/3 = n(CCl2F2)/3

n(CCl4) = n(CCl2F2)

n(CCl2F2) = n(CCl4) = 0.001300 mol

(0.001300 mol CCl2F2) × (120.91 g CCl2F2 / 1 mol CCl2F2) = 0.157183 g CCl2F2 = 0.157 g CCl2F2


Answer: 0.157 grams of CCl2F2 is formed.